The Bok Chicken Factory is trying to figure out how to minimize the cost of producing 1200 units of chicken parts. The production function is q = 100L0.5 K0.5. The wage rate is birr 9 per hour and the rental rate on capital is birr 4 per machine hour.



Find the minimum cost of producing 1200 units.



Find the maximum output that can be produced for a total cost of birr 720.

1
Expert's answer
2022-01-24T11:02:43-0500

w=9

r=4

q=1200

"q=100L^\frac{1}{2}K^\frac{1}{2}"

Derive MPL and MPK


MPL = "\frac{\partial Q} {\partial L}" = 50L-0.5K0.5


MPK = "\frac{\partial Q} {\partial K}" = 50L0.5K-0.5


"\frac{MPL}{MPK}" = "\frac{w}{r}"

w = 9

r = 4

50L-0.5K0.5 "\div" 50L0.5K-0.5 = "\frac{9}{4}"


"\frac {50L^{-0.5}K^{0.5}}{50L^{0.5}K^{-0.5}}=\frac{9}{4}"


"\frac{K}{L}=\frac{9}{4}"


K="\frac{9}4L"


Substitute this value in the production function

"q=100(\frac{9}{4}L)^{0.5}L^{0.5}"


"1200=100(\frac{13}{4}L)^{0.5}"


"L^{0.5}=\frac{1200}{(100)(3.25)^{0.5}}"


"L=(\frac{1200}{(100)(3.25^{0.5}})^2"

"= \frac{1440000}{32500}= 44.31"


"K= \frac{9}{4}L"

"K= \frac{9}{4}\times 44.31= 99.69"


TC= wL+rK

720=9L+4K


But L="\frac {9}{4}K"

"720= 9L+\frac{9}{4}L"

"720=\frac{45}{4}L"


"L=\frac{(720\times 4)}{45}= 64"


"K=\frac{9}{4}\times 64= 144"


Total Output=144+64= 208